OpenStudy (anonymous):

For element Cl a) Calculate the number of moles in 9.36g b) Calculate the number of grams of 0.104 mol c) How many moles are there in 9.56x10^22 atoms of the element?

OpenStudy (cuanchi):

You need to know the atomic mass of the Cl element. You can find it in a periodic table a) you can have a conversion factor 1 mol Cl = atomic mass of Cl and multiply your 9.36g by the conversion factor. b) Whith the inverse conversion factor from a you can calculate the g in 0.104 mol c) you will need the Avogadro's number (6.02 x 10^23) to calculate this value

OpenStudy (anonymous):

for a i got 0.246 does this seem correct @cu

OpenStudy (anonymous):

@Cuanchi

OpenStudy (abb0t):

This type of thought process is considered dimensional analysis. Here is how you set it up, \(\sf always\): \(\sf \color{red}{given} \times \frac{units~you~want}{units~of~given} = \color{blue}{want}\)

OpenStudy (abb0t):

For a, you are given 9.36 grams so 9.36 grams is what you're given. Now, look at your periodic table. Remeber the units on the periodic table are given in grams PER mole, which means: \(\sf \color{red}{\frac{grams}{mole}}\) so if you want \(\sf moles\) you should FLIP the units. Which means, you'll be dividing by the grams you find on the periodic table. \(\sf grams~of~Cl \times \frac{moles}{grams}\) notice units of grams will cancel it. it is the same as algebra, as if you had \(\sf a \times \frac{b}{a} = b\)

OpenStudy (anonymous):

I got a i need help with b @abb

OpenStudy (anonymous):

@abb0t

OpenStudy (abb0t):

You're basically doing the opposite. But same process. Still focusing on the element Chlorine.

OpenStudy (abb0t):

Refer to your periodic table. If you don't have one, use this: http://www.ptable.com/

OpenStudy (cuanchi):

a) 9.36/35.45= 0.264 mol